OpenStudy (anonymous):

If the velocity of the object is linear, then the acceleration is v(t)=, and the velocity function is v(t)=.

OpenStudy (anonymous):

v prime(t)= 1ft/s^2

OpenStudy (amistre64):

linear as in constant? or as in v(t) = at ?

OpenStudy (anonymous):

what is the v prime function (t)

OpenStudy (amistre64):

i cant make sense of what you are asking ....

OpenStudy (anonymous):

how do i write a velocity function?

OpenStudy (amistre64):

acceleration is the change in velocity, if the velocity is not changing, the acceleration is 0 you write a velocity function by integrating acceleration

OpenStudy (amistre64):

if: a(t) = 3x; then v(t) = (3x^2)/2 + C

OpenStudy (amistre64):

if: a(t) = 0; then v(t) = C

OpenStudy (anonymous):

so if v prime(t)= 1ft/s^2 then what would v(t)=?

OpenStudy (amistre64):

lol, the words threw me off :) a(t) = v'(t) since a(t) = 1; then v(t) = \(\int 1~ dt\)

OpenStudy (anonymous):

so what is the velocity function?

OpenStudy (amistre64):

ok, since my way is apparently foreign to you; what methods do you know of that would get you there?

OpenStudy (anonymous):

i know what the a(t) is, but how do i write that into a function of the velocity?

OpenStudy (amistre64):

by integrating a(t); that is one method that uses calculus

OpenStudy (anonymous):

how do i write the formula then?

OpenStudy (amistre64):

\[a(t)=1\]\[\int a(t)~dt=\int 1~dt\]

OpenStudy (anonymous):

that is the whole question.. i just need help on the last one..

OpenStudy (anonymous):

Without calculus:\[acceleration=v'(t)=a=\frac{\text{ change of velocity}}{time}\]If the velocity is constant, there is no change, if it is linear, there is a constant change.

OpenStudy (anonymous):

so what is the last part?